# RPRO: Ranked Preference Reinforcement Optimization for Enhancing Medical QA and Diagnostic Reasoning

### v1

@@ 1 Introduction
Recent advances in large language models (LLMs) signal significant promise for medical applications, especially in clinical reasoning and diagnostic support systems [1, 2, 3, 4]. However, deploying AI in medicine requires exceptional reliability, making factual accuracy and thorough coverage of the data paramount. Medical reasoning introduces challenges distinct from those in general domains: clinical decisions require a comprehensive consideration of relevant factors, strict factual accuracy, and the avoidance of redundant information that could obscure critical insights.
Current approaches to medical reasoning primarily rely on chain-of-thought (CoT) prompting techniques [5, 6], which have shown effectiveness in improving reasoning capabilities across various domains. However, these methods typically generate single reasoning chains without systematic quality assessment or iterative refinement mechanisms. Moreover, existing evaluation frameworks often rely on overly simplistic binary metrics that fail to capture the nuanced quality required for medical reasoning or require extensive human annotation, limiting their suitability for automated quality control where domain expertise is essential.
Recent work in preference learning and reinforcement learning from human feedback (RLHF) [7, 8] shows promise in aligning AI systems with human preferences. Direct Preference Optimization (DPO) [9] has shown promising results by learning from preference data. However, these methods typically rely on pairwise comparisons, which fail to capture all the ranking information in human preference datasets. Many real-world scenarios involve humans ranking multiple candidates, not just providing binary choices. For example, when they evaluate reasoning chains, mathematical solutions, or creative content, evaluators assign rankings that reflect nuanced quality differences. Reducing this rich data to pairwise comparisons can lose valuable training signals and lead to suboptimal solutions.
@@ 2.1 Medical LLMs: QA, Diagnosis, and EHR
Large language models (LLMs) have been used for medical tasks like knowledge-based question answering, diagnostic reasoning on clinical vignettes, and patient-specific reasoning over electronic health records (EHR) [10, 11, 12, 13]. Domain-structured prompts and chain-of-thought (CoT) approaches consistently improve auditability and reduce unsupported claims by clearly separating decomposition and background from justification or diagnosis [14, 15, 16]. In patient-specific contexts, retrieval-augmented generation (RAG) and citation-style evidence tracing help ensure faithfulness to chart data and guidelines [17, 12, 13]. Recent work extends beyond answer accuracy, emphasizing multidimensional evaluation such as coverage of relevant findings and alignment with context, thereby better matching clinical expectations for completeness and correctness. In this work, we use domain-specific CoT templates (QA and diagnosis), automatically score rationales on coverage, factuality, and redundancy with an LLM judge, and make light score-guided revisions.
@@ 2.2 Reinforcement-Learned Medical LLMs
Reinforcement learning from human feedback (RLHF) and related preference-based methods have become central to aligning LLM behavior beyond supervised fine-tuning [8, 18, 7]. In reasoning-heavy domains, variants that regularize to a reference policy while optimizing rewards derived from preference ordering or task-specific outcomes have shown improved stability and faithfulness.
In the medical domain, reinforcement-learned LLMs have been applied to tasks such as clinical question answering and diagnostic reasoning [19, 20]. These methods typically employ reinforcement learning techniques such as policy optimization with preference feedback [8, 7] to guide models toward multi-step clinical reasoning, integration of biomedical knowledge, and reliable elimination of alternative diagnoses. Nevertheless, prior studies also emphasize persistent challenges in maintaining stability and factual grounding in high-stakes clinical scenarios, indicating that preference-optimized training plays an important role in reducing hallucination risks [19].
@@ 2.3 Ranking-based Optimization in Clinical Reasoning
Recent studies have expanded preference alignment for LLMs from pairwise to listwise settings. Most methods focus on pairwise comparisons. Yet, listwise approaches have shown effectiveness in other domains [21, 22, 23]. For example, ALMupQA by [24] integrates multi-perspective ranking alignment for code QA. IRPO, introduced by [25], is a framework that incorporates graded relevance and positional importance. These works highlight the impact of ranking-based feedback on aligning LLM outputs with diverse user expectations. However, medical multiple-choice QA often needs more sophisticated explanatory discourse and clinical reasoning. In this work, we evaluate our framework on multiple medical reasoning benchmarks. We use a probabilistic refinement mechanism, generate preference data with RPRO, and combine linear rewards for ranked candidate optimization.
@@ 3.4 Medical Prompt Instruction and Reasoning
Previous research shows that effective prompt template instructions can improve model performance on CoT reasoning [26, 5, 6]. We use structured reasoning templates based on established clinical and biomedical reasoning patterns. Table C presents a general medical QA template that applies a four-step reasoning process, using binary classification at each step according to the validation criteria from Section 3.3. Specifically, background knowledge retrieval gathers relevant biomedical facts and principles from medical literature. Logical connection links the retrieved knowledge to the question context. The justification step synthesizes the reasoning to support the conclusion.
@@ 4.1 Ranking Preference Optimization Algorithm
where $|c_{j}|$ denotes the sequence length of the CoT, and $c_{j,t}$ represents the token at position $t$ in the reasoning chain. Following the Bradley-Terry model [27], the probability that CoT sequence $c_{i}$ is preferred over sequence $c_{j}$ is:
@@ 4.2 Linear Rewards From Rank
GRPO is one of the widely used algorithms that employs entropy regularization to stabilize policy gradient–based learning processes. However, it still exhibits sensitivity to parameter tuning, which can lead to unstable convergence performance or even divergence during training [28, 29, 30]. To translate the complete ranking of candidate reasoning chains into effective training signals, we adopt a linear reward shaping scheme. We then define the advantage value $a_{j}$ of candidate $c_{j}$ as:
@@ 5.1 Experimental Setting
Datasets. We evaluate our approach on two benchmark medical QA datasets. PubMedQA [10] contains biomedical research questions derived from PubMed abstracts, requiring yes/no/maybe answers based on scientific evidence. MedQA-USMLE [31] consists of multiple-choice diagnostic questions from the United States Medical Licensing Examination, testing clinical reasoning and diagnostic capabilities. These datasets represent complementary aspects of medical reasoning: research-oriented knowledge synthesis and clinical diagnostic decision-making.
@@ 5.2 Baselines
We compare against several established language models across different scales and medical specializations. General-purpose models include Gemma 2B and 7B [32], and Mistral 7B [33]. Medical-specialized models include MedAlpaca 7B and 13B [34], fine-tuned specifically for medical tasks, and BioMistral 7B [35], trained on biomedical corpora. These baselines represent the current state-of-the-art in both general and domain-specific language modeling for medical applications.
@@ 5.5 Quantitative Results
Reasoning Results from PubMedQA and MedQA-USMLE Tasks. The reasoning examples shown in Table C in Appendix A.2 provide representative cases generated by our model using the proposed prompt templates. These examples demonstrate how our approach enables structured, interpretable, and domain-specific reasoning across both biomedical and clinical question answering tasks.
@@ A.1 Prompt Templates
Prompt Template (PubMedQA style) This is your content inside the tcolorbox. Make sure you don’t break the syntax.
@@ A.2 Model-Generated Reasoning Cases
The following reasoning cases, shown in Table C, present representative outputs generated by our model using the proposed prompt templates. These cases demonstrate how our approach enables structured, interpretable, and domain-specific reasoning across both biomedical and clinical question answering tasks.
@@ Appendix C Dataset Format (JSONL Example)
1{
2 "id": "PMQA_001",
3 "question": "Can patient coaching reduce disparities?",
4 "context": "This study evaluated whether structured patient coaching improves adherence and reduces disparities ...",
5 "answer": "Yes",
6 "ranked": [
@@
9 ],
10 "scores": [
11 0.90,
12 0.65
13 ],
14 "breakdowns": [
15 {"coverage": 4.5, "factual_accuracy": 4.7, "redundancy": 0.2},
16 {"coverage": 3.5, "factual_accuracy": 3.2, "redundancy": 0.7}
17 ]
18}

### v2

# The Second Workshop on GenAI for Health: Potential, Trust, and Policy Compliance

@@ 1 Introduction
Recent advances in large language models (LLMs) signal significant promise for medical applications, especially in clinical reasoning and diagnostic support systems singhal2023large; thirunavukarasu2023large. However, deploying AI in medicine requires exceptional reliability, making factual accuracy and thorough coverage of the data paramount. Medical reasoning introduces challenges distinct from those in general domains: clinical decisions require a comprehensive consideration of relevant factors, strict factual accuracy, and the avoidance of redundant information that could obscure critical insights.
Current approaches to medical reasoning primarily rely on chain-of-thought (CoT) prompting techniques wei2022chain; kojima2022large, which have shown effectiveness in improving reasoning capabilities across various domains. However, these methods typically generate single reasoning chains without systematic quality assessment or iterative refinement mechanisms. Moreover, existing evaluation frameworks often rely on overly simplistic binary metrics that fail to capture the nuanced quality required for medical reasoning or require extensive human annotation, limiting their suitability for automated quality control where domain expertise is essential.
Recent work in preference learning and reinforcement learning from human feedback (RLHF) ouyang2022training; christiano2017deep shows promise in aligning AI systems with human preferences. Direct Preference Optimization (DPO) rafailov2023direct has shown promising results by learning from preference data. However, these methods typically rely on pairwise comparisons, which fail to capture all the ranking information in human preference datasets. Many real-world scenarios involve humans ranking multiple candidates, not just providing binary choices. For example, when they evaluate reasoning chains, mathematical solutions, or creative content, evaluators assign rankings that reflect nuanced quality differences. Reducing this rich data to pairwise comparisons can lose valuable training signals and lead to suboptimal solutions.
@@ 2.1 Medical LLMs: QA, Diagnosis, and EHR
Large language models (LLMs) have been used for medical tasks like knowledge-based question answering, diagnostic reasoning on clinical vignettes, and patient-specific reasoning over electronic health records (EHR) (jin2019pubmedqa; jin2021medqa; shi2024ehragent; zakka2024almanac). Domain-structured prompts and chain-of-thought (CoT) approaches consistently improve auditability and reduce unsupported claims by clearly separating decomposition and background from justification or diagnosis (wei2022cot; singhal2023medpalm; singhal2023medpalm2). In patient-specific contexts, retrieval-augmented generation (RAG) and citation-style evidence tracing help ensure faithfulness to chart data and guidelines lewis2020rag; shi2024ehragent; zakka2024almanac. Recent work extends beyond answer accuracy, emphasizing multidimensional evaluation such as coverage of relevant findings and alignment with context, thereby better matching clinical expectations for completeness and correctness. In this work, we use domain-specific CoT templates (QA and diagnosis), automatically score rationales on coverage, factuality, and redundancy with an LLM judge, and make light score-guided revisions.
@@ 2.2 Reinforcement-Learned Medical LLMs
Reinforcement learning from human feedback (RLHF) and related preference-based methods have become central to aligning LLM behavior beyond supervised fine-tuning christiano2017deep; ziegler2019fine; ouyang2022training. In reasoning-heavy domains, variants that regularize to a reference policy while optimizing rewards derived from preference ordering or task-specific outcomes have shown improved stability and faithfulness.
In the medical domain, reinforcement-learned LLMs have been applied to tasks such as clinical question answering and diagnostic reasoning singhal2023towards; nori2023capabilities. These methods typically employ reinforcement learning techniques such as policy optimization with preference feedback christiano2017deep; ouyang2022training to guide models toward multi-step clinical reasoning, integration of biomedical knowledge, and reliable elimination of alternative diagnoses. Nevertheless, prior studies also emphasize persistent challenges in maintaining stability and factual grounding in high-stakes clinical scenarios, indicating that preference-optimized training plays an important role in reducing hallucination risks singhal2023towards.
@@ 2.3 Ranking-based Optimization in Clinical Reasoning
Recent studies have expanded preference alignment for LLMs from pairwise to listwise settings. Most methods focus on pairwise comparisons. Yet, listwise approaches have shown effectiveness in other domains liu2024lipo; pesaran2025lpoi; cai2025kpo. For example, ALMupQA by yang2024aligning integrates multi-perspective ranking alignment for code QA. IRPO, introduced by wu2025context, is a framework that incorporates graded relevance and positional importance. These works highlight the impact of ranking-based feedback on aligning LLM outputs with diverse user expectations. However, medical multiple-choice QA often needs more sophisticated explanatory discourse and clinical reasoning. In this work, we evaluate our framework on multiple medical reasoning benchmarks. We use a probabilistic refinement mechanism, generate preference data with RPRO, and combine linear rewards for ranked candidate optimization.
@@ 3.4 Medical Prompt Instruction and Reasoning
Previous research shows that effective prompt template instructions can improve model performance on CoT reasoning zhu2024llms; wei2022chain; kojima2022large. We use structured reasoning templates based on established clinical and biomedical reasoning patterns. Table 3 presents a general medical QA template that applies a four-step reasoning process, using binary classification at each step according to the validation criteria from Section 3.3. Specifically, background knowledge retrieval gathers relevant biomedical facts and principles from medical literature. Logical connection links the retrieved knowledge to the question context. The justification step synthesizes the reasoning to support the conclusion.
@@ 4.1 Ranking Preference Optimization Algorithm
where $|c_{j}|$ denotes the sequence length of the CoT, and $c_{j,t}$ represents the token at position $t$ in the reasoning chain. Following the Bradley-Terry model bradley1952rank, the probability that CoT sequence $c_{i}$ is preferred over sequence $c_{j}$ is:
@@ 4.2 Linear Rewards From Rank
GRPO is one of the widely used algorithms that employs entropy regularization to stabilize policy gradient–based learning processes. However, it still exhibits sensitivity to parameter tuning, which can lead to unstable convergence performance or even divergence during training leahy2022convergence; schulman2017proximal; mnih2016asynchronous. To translate the complete ranking of candidate reasoning chains into effective training signals, we adopt a linear reward shaping scheme. We then define the advantage value $a_{j}$ of candidate $c_{j}$ as:
@@ 5.1 Experimental Setting
Datasets. We evaluate our approach on two benchmark medical QA datasets. PubMedQA jin2019pubmedqa contains biomedical research questions derived from PubMed abstracts, requiring yes/no/maybe answers based on scientific evidence. MedQA-USMLE jin2021disease consists of multiple-choice diagnostic questions from the United States Medical Licensing Examination, testing clinical reasoning and diagnostic capabilities. These datasets represent complementary aspects of medical reasoning: research-oriented knowledge synthesis and clinical diagnostic decision-making.
@@ 5.2 Baselines
We compare against several established language models across different scales and medical specializations. General-purpose models include Gemma 2B and 7B mesnard2024gemma, and Mistral 7B jiang2023mistral. Medical-specialized models include MedAlpaca 7B and 13B han2023medalpaca, fine-tuned specifically for medical tasks, and BioMistral 7B labrak2024biomistral, trained on biomedical corpora. These baselines represent the current state-of-the-art in both general and domain-specific language modeling for medical applications.
@@ 5.5 Quantitative Results
Reasoning Results from PubMedQA and MedQA-USMLE Tasks. The reasoning examples shown in Table 3 in Appendix A.2 provide representative cases generated by our model using the proposed prompt templates. These examples demonstrate how our approach enables structured, interpretable, and domain-specific reasoning across both biomedical and clinical question answering tasks.
@@ 6 Conclusion and Future Work
In this paper, we present a novel approach for enhancing medical reasoning by combining ranked preference reinforcement optimization with probabilistic reasoning refinement. Our method introduces task-adaptive Chain-of-Thought generation, employs domain-specific reasoning templates, and incorporates a probabilistic quality assessment framework that enables targeted self-reflection and refinement. We evaluate our 1.1B parameter model on PubMedQA and MedQA-USMLE, demonstrating substantial improvements over much larger baseline models. These results validate that quality-driven reasoning enhancement is more effective than simple parameter scaling. Ablation studies further highlight the critical roles of reasoning refinement and ranked preference optimization.
@@ A.1 Prompt Templates
Prompt Template (PubMedQA style)
@@ A.2 Model-Generated Reasoning Cases
The following reasoning cases, shown in Table 3, present representative outputs generated by our model using the proposed prompt templates. These cases demonstrate how our approach enables structured, interpretable, and domain-specific reasoning across both biomedical and clinical question answering tasks.
@@ Appendix C Dataset Format (JSONL Example)
"id": "PMQA_001",
"question": "Can patient coaching reduce disparities?",
"context": "This study evaluated whether structured patient coaching improves adherence and reduces disparities ...",
"answer": "Yes",
"ranked": [
@@
"scores": [
0.90,
0.65,
"breakdowns": [
{"coverage": 4.5, "factual_accuracy": 4.7, "redundancy": 0.2},
{"coverage": 3.5, "factual_accuracy": 3.2, "redundancy": 0.7},
Table 3: Reasoning results for PubMedQA and MedQA-USMLE tasks.